OpenStudy (anonymous):

Evaluate 10 + 4(3 + 2) + 5 + (12 ÷ 6).

OpenStudy (anonymous):

37

OpenStudy (anonymous):

37

OpenStudy (anonymous):

10 + 4*3 + 4*2 + 5 + 2 = 10 + 12 + 8 +7 = 22 + 8 +7 = 30 + 7 =37